OpenStudy (anonymous):

Please explain me how to get the answer (4/1)^2 divided by (4/1 - 1/3)

geerky42 (geerky42):

You mean this? \[\dfrac{\left(\dfrac{4}{1}\right)^2}{\dfrac{4}{1}-\dfrac{1}{3}}\]

OpenStudy (anonymous):

12/11

OpenStudy (anonymous):

if u need help with such questions visit mathway.com

OpenStudy (anonymous):

\[\left( \frac{ 4 }{ 1} \right)^2 \div \left( \frac{ 4 }{ 1} - \frac{ 1 }{ 3}\right)\]

OpenStudy (anonymous):

12/11 is wrong @aballer101\

OpenStudy (anonymous):

thats wat the website said. u can also try webmath.com

OpenStudy (anonymous):

@geerky42 could you help us out?

OpenStudy (anonymous):

Well @aballer101 my school work on mymathlab says its incorrect...i just need an understand on how to solve...i am not looking for an answer

OpenStudy (anonymous):

@robtobey can you help me